OpenStudy (anonymous):

12% of all drivers do not have a valid drivers license, 6% of all drivers have no insurance, and 4% have neither what is the probability that a randomly selected driver either fails to have a valid license or fails to have insurance is about a) 0.18 b) 0.2 c) 0.22 d) 0.072 e) 0.14

OpenStudy (xapproachesinfinity):

you are looking for \[P(A\cup B)=P(A)+P(B)-P(A\cap B)\] A event that a driver fails to have a valid licence B event that a driver fails to have insurance so \[P(A)=\frac{12}{100}=0.12\\ P(B)=\frac{6}{100}=0.06\\ P(A\cap B)=0.12\times 0.06=0.0072\] do the rest...

OpenStudy (xapproachesinfinity):

perhaps i assumed not independent ?!

OpenStudy (mathmate):

\(P(A\cap B)\) is given to be 0.04.

OpenStudy (xapproachesinfinity):

yea i just realized

OpenStudy (xapproachesinfinity):

i get 0.176 as the prob

OpenStudy (xapproachesinfinity):

did they round to 0.18

OpenStudy (xapproachesinfinity):

or did i assume wrong situation

OpenStudy (mathmate):

It's 0.04, and not 0.004, you were almost there.

OpenStudy (xapproachesinfinity):

oh then my bad heheh i don't know where did i add one zero lol

OpenStudy (xapproachesinfinity):

got 0.14

OpenStudy (xapproachesinfinity):

made me think that i assumed wrong situation lol

OpenStudy (mathmate):

Your original formula is still correct. I usually let the OP to work out the numbers, otherwise she doesn't get anything out of the exercise.

OpenStudy (xapproachesinfinity):

thanks for the correction yeah my intention was to let her work it out but then got this error lol

OpenStudy (mathmate):

No problem, I know of the complications! :)

OpenStudy (xapproachesinfinity):

:)

OpenStudy (xapproachesinfinity):

so @trice2964 did you get this

OpenStudy (anonymous):

yes! thank you! so basically you add up all the percentages, 12%+6%+4% which equals 0.22 and then you divide 0.04/.22 and the answer is about .18? @xapproachesinfinity

OpenStudy (xapproachesinfinity):

oh no i did 12% + 6% - 4% which give me 0.14 not 0.18

OpenStudy (xapproachesinfinity):

see the first formula that i wrote

OpenStudy (anonymous):

yes but why do you suntract 4% instead of adding it? @xapproachesinfinity

OpenStudy (xapproachesinfinity):

because if we have two events and we want to know the probability of either one occurring we add the probability of each event but if we did that we over counted so we subtract the probability of both happing at the same time this for the independent events \[P(A\cup B)=P(A)+P(B)-B(A \cap B)\] that exactly what this formula says

OpenStudy (xapproachesinfinity):

and we said that \[P(A\cap B)=0.04\]

OpenStudy (xapproachesinfinity):

i meant happening not happing lol

OpenStudy (xapproachesinfinity):

perhaps van diagram can help visualize this hehe

OpenStudy (mathmate):



OpenStudy (xapproachesinfinity):

exactly :)

OpenStudy (anonymous):

wait where did the 2% and 8% percent come from?! I'm sorry you guys don't think I'm a idiot lol. I've been struggling with probability really badly, I think i think to hard about it because what I originally did was just add 0.12+0.06+0.04=.22 and then I did 0.04/.22 which equals 0.18. But I see where I was wrong now kind of lol

OpenStudy (xapproachesinfinity):

so we have 12% chance for a driver with n valid licence also 6% chance that the driver has no insurance lastly, the chance a drives does not have both insurance and a valid license is 4% 12%-4%=8% 6%-4%=2% 4% is the a shared percentage

OpenStudy (xapproachesinfinity):

the more you struggle and the efforts you out the more you understand :) keep it up

OpenStudy (anonymous):

Ohhhh I understand it a little bit better now, you subtract the 4% because its shared? @xapproachesinfinity

OpenStudy (xapproachesinfinity):

yes exactly

OpenStudy (xapproachesinfinity):

if we didn't do we over counted then

OpenStudy (anonymous):

and if it's not shared you would add it? @xapproachesinfinity

OpenStudy (xapproachesinfinity):

if there is no shared parts then you add straight depends on the problem

OpenStudy (xapproachesinfinity):

that happen in mutually exclusive events

OpenStudy (xapproachesinfinity):

see here http://www.mathsisfun.com/data/probability-events-mutually-exclusive.html

OpenStudy (anonymous):

okay! well let me try this and tell me if I it's correct, but not the right answer because I want to try and get it myself lol. But Its the same scenario -The probability that a randomly selected driver fails to have insurance, given that he fails to have a valid license is about a)0.67 b)0.33 c) 0.06 d)0.50 e).01 Would if be 0.06? because I subtracted the 12% from the 6% @xapproachesinfinity

OpenStudy (xapproachesinfinity):

oh no this one is different

OpenStudy (xapproachesinfinity):

this conditional probability

OpenStudy (xapproachesinfinity):

this has the formula \[P(A/B)=\frac{P(A\cap B)}{P(B)}\]

OpenStudy (xapproachesinfinity):

conditional probabilities are different lol

OpenStudy (xapproachesinfinity):

let A be event that a driver fails to have insurance B be event that a driver fails to have a valid license

OpenStudy (anonymous):

okay I'm going to work the problem out now

OpenStudy (xapproachesinfinity):

okay!

OpenStudy (anonymous):

@xapproachesinfinity okay i think i have it now! lol So since it is conditional probability which means it can be affected by previous effects! Using the formula I divided 0.06 by 0.12 and I got d 0.50!

OpenStudy (xapproachesinfinity):

should be 0.04/0.12

OpenStudy (anonymous):

oh my gosh . i give up =((((

OpenStudy (xapproachesinfinity):

remember that shared part that is what goes into numerator

OpenStudy (anonymous):

oh wait you use 0.04 because the person doesnt have drivers licenses or insurance!

OpenStudy (xapproachesinfinity):

oh trying i better than nothing and none of got it just like that heheh constant practice....

OpenStudy (xapproachesinfinity):

no P(a and b) means a driver doesn't have good license and insurance both at one time

OpenStudy (xapproachesinfinity):

watch some videos perhaps that would help you youtube has some good videos

OpenStudy (xapproachesinfinity):

I still get stuck on probability heheh so don't get frustrated lol

OpenStudy (anonymous):

Maybe so! I've never been good at word problems though lol. But thank you for helping me!!!

OpenStudy (xapproachesinfinity):

no problem probability is a fun subject though lol just practice, we all start at nothing but we continue to do more and more problems and that comes to be helpful :)

OpenStudy (xapproachesinfinity):

to work a word problem you need a strategy 1) read effectively the problem ( 2) understand what is asked 3) think about what you know 4) make connection 5) start attacking the problem 6) solve and review

OpenStudy (xapproachesinfinity):

i'm sure you can find a good strategy that is much better :)

OpenStudy (xapproachesinfinity):

organizing the ideas is key to problem solving

OpenStudy (anonymous):

I get confused because I don't know what formula to use for each problem and I get confused on what to put into the formula!

OpenStudy (xapproachesinfinity):

yeah i understand you :) review the notes of the lecture before doing problems of course see a lot of example of how to use the formulas after that do problems

OpenStudy (anonymous):

okay for the last problem! why would you divide 0.06 by 0.12 because in the problem it says the probability of a driver who fails to have insurance, and a driver who fails to have a drivers licenses, so those percentages would be 6% and 12%

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

first its not 0.06/0.12 it is 0.04/0.12 and that's because your question said a driver fails to have insurance , "given" that a driver fails to have valid license

OpenStudy (xapproachesinfinity):

highlight "given"

OpenStudy (xapproachesinfinity):

we are looking for the probability that a driver doesn't have insurance given that we know he does not have a valid license

OpenStudy (xapproachesinfinity):

and that's we define as \[P(A/B)=\frac{P(A\cap B)}{P(B)}\]

OpenStudy (xapproachesinfinity):

we know that \[P(A\cap B)=0.04\]

OpenStudy (mathmate):

 "Given" means it has already happened. "So given he fails to have valid licence" means you choose those that do not have valid insurance (4%) (event A \( \cap \) B) \(out~ of\) the 12% (event B)